How many Na+ ions are contained in 4.62g Na3PO4?


1
Expert's answer
2022-02-16T05:55:05-0500

Solution:

The molar mass of Na3POis 163.94 g/mol

Therefore,

Moles of Na3PO4 = (4.62 g Na3PO4) × (1 mol Na3PO4 / 163.94 g Na3PO4) = 0.02818 mol Na3PO4


Na3PO4 → 3Na+ + PO43-

Therefore,

1 mol of Na3PO4 contains 3 mol of Na+ ions

Thus, 0.02818 mol of Na3PO4 contains:

(0.02818 mol Na3PO4) × (3 mol Na+ ions / 1 mol Na3PO4) = 0.08454 mol Na+ ions


One mole of any substance contains 6.022×1023 particles.

Therefore,

Number of Na+ ions = (0.08454 mol) × (6.022×1023 ions / 1 mol) = 5.09×1022 ions


Answer: 5.09×1022 of Na+ ions